OpenStudy (ray_sihota):

having trouble with putting these equations into y=mx+b format write equation of a line passing through (-1,4) slope -1/3 the answer in the back of the book is -1/3+11/3 i do not understand how they got 11/3

OpenStudy (agreene):

so, i would start by putting it in point-slope form, since they gave you a point: \[y-y_1=m(x-x_1)\] \[y-4=\frac{-1}{3}(x-1)\]

OpenStudy (agreene):

then solve that for y=mx+b form

OpenStudy (agreene):

i messed up the sign there it should be y-4= -1/3(x+1) not x-1

OpenStudy (ray_sihota):

got it thank you